OpenStudy (anonymous):

If g(x)=-x^2+6x-1 . What's g(-2) ?

OpenStudy (shaik0124):

-17 IS THE ANSWER

OpenStudy (anonymous):

Show Work Please . I'm Just Not Understanding This One .

OpenStudy (shaik0124):

-(-2^2)+(6*(-2)-1=-4-12-1=-17

OpenStudy (shaik0124):

I HAVE SOLVED FIRST

OpenStudy (shaik0124):

HOW 2 GET ACHIEVEMENTS

OpenStudy (shaik0124):

RPLY ANYONE

OpenStudy (anonymous):

Oh , I'm Not Sure . Sorry .

OpenStudy (shaik0124):

BUT U HAVE 5 ACHIEVEMENTS ON UR PROFILE

OpenStudy (anonymous):

Oh , Well I Don't Know How I Got Them .